OpenStudy (anonymous):

At what point on the curve y = 5 + 2e^x − 4x is the tangent line parallel to the line 4x − y = 2?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

do you know what the derivative of \(y=5+2e^x-4x\) is? you need that first

OpenStudy (anonymous):

2e^x-4x? @satellite73

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Help please^^

OpenStudy (triciaal):

parallel lines have the same slope.

OpenStudy (danjs):

the derivative is fancy for the slope of a tangent line at a point on a curve, compare that to the slope of that line

OpenStudy (triciaal):

rephrasing the question when is the first derivative of the curve the same value as the slope of the line given?

OpenStudy (triciaal):

do you know how to continue now?

OpenStudy (anonymous):

I am still confused, what do I find the slope of ? @triciaal

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

hi

OpenStudy (anonymous):

I dont get this question

OpenStudy (danjs):

They give you some curvy function y = f(x) = 5 + 2e^x - 4x and a line (y = mx+b) 4x - y = 2 rearrange y = 4x - 2 slope m=4

OpenStudy (danjs):

They want the point (x,y) on that curvy f(x), where the tangent line has same slope m=4 of that other given line

OpenStudy (danjs):

got that so far?

OpenStudy (danjs):